# Some Aspects of Three-Quark Potentials (Part II)

### v2

@@ II Preliminaries
Figure 1: A typical baryonic configuration in five dimensions. The heavy quarks are denoted by $Q$, the baryon vertex by $V$, and its projection onto the boundary by $Y$. The strings are labeled by 1, 2, 31231,\,2,\,31 , 2 , 3. The forces acting on the vertex are indicated by the arrows.
@@
[…]
@@ III Isosceles Triangle Geometry
which are valid for $20\leq\beta\leq(\pi-\theta)/$ 2.
Due to axial symmetry, the force balance equation (2.5) reduces to two non-trivial components. Those are the $y$- and $r$-components:
@@
Here $\alpha_{i}$ is the tangent angle of the i-string (see Figure 13), and $v=\mathsf{s}r_{v}^{2}$, with $r_{v}$ being the position of the vertex along the $r$ axis. The coefficient $\mathsf{k}$ is defined below in Eq.(3.4). By symmetry, we have $\alpha_{1}=\alpha_{2}$. It is important to note that for $\beta>(\pi-\theta)/2$, the string configuration doesn’t exist, as the force balance equation has no solution in this case.
@@ III.1.1 Configuration I
where the functions ${\cal L}^{\pm}$ are defined in Appendix A and $\lambda_{3}=-\text{ProductLog}(-v\,\text{e}^{-v}/\cos\alpha_{3})$, as follows from Eq.(B.15). $\ell$ can be extracted from Eq.(3.1) by substituting $|Q_{1}Y|={\cal L}^{+}(\alpha_{1},v)/\sqrt{\mathsf{s}}$, and the energy is obtained from Eqs.(2.3), (B.11), and (B.17). The result is [^4] The superscript refers to a string configuration. We often omit it to reduce clutter.
@@
Here $\mathsf{g}=\frac{R^{2}}{2\pi\alpha^{\prime}}$, $\mathsf{k}=\frac{\tau_{v}}{3\mathsf{g}}$, and $c$ is a normalization constant. Combined with the geometrical constraint and the force balance equations, this yields the function $E^{\,\text{\tiny I}}(\ell)$ in parametric form. [^5] The allowed range of $v$ is discussed in subsection B and Appendix D.
Now consider the $-\ell$ behavior of $E^{\,\text{\tiny I}}$. This limit corresponds to $v\rightarrow 0$ as follows from the asymptotic behavior of the function ${\cal L}^{+}$. In this case, the force balance equations and the geometrical constraint reduce to
@@
$\cos\bigl{(}\beta+\tfrac{\theta}{2}\bigr{)}\sqrt{\cos\alpha_{3}}\,I(\cos^{2}\alpha_{1},\tfrac{3}{4},\tfrac{1}{2})=\sin\tfrac{\theta}{2}\sqrt{\cos\alpha_{1}}\bigl{(}1+I(\sin^{2}\alpha_{3},\tfrac{1}{2},\tfrac{3}{4})\bigr{)}\,,$ (3.6)
@@ III.1.2 Configuration II
This configuration is governed by the same action as before. As a result, the force balance equations (3.2) remain valid. In fact, since $\alpha_{1}$ now changes the sign from positive to negative, the corresponding expressions for the geometrical constraint, leg length and energy can be obtained simply by replacing ${\cal L}^{+}$ and ${\cal E}^{+}$ with ${\cal L}^{-}$ and ${\cal E}^{-}$. Thus, we immediately arrive at
@@
It is now straightforward to analyze the $-\ell$ behavior of $E^{\,\text{\tiny II}}(\ell)$. The force balance equations (3.5) apply in this case as well. A similar analysis to the one carried out previously shows that the geometrical constraint becomes
$\cos\bigl{(}\beta+\tfrac{\theta}{2}\bigr{)}\sqrt{\cos\alpha_{3}}\bigl{(}1+I(\sin^{2}\alpha_{1},\tfrac{1}{2},\tfrac{3}{4})\bigr{)}=\sin\tfrac{\theta}{2}\sqrt{\cos\alpha_{1}}\bigl{(}1+I(\sin^{2}\alpha_{3},\tfrac{1}{2},\tfrac{3}{4})\bigr{)}\,.$ (3.10)
@@
Now we would like to examine the $-\ell$ behavior of $E^{\,\text{\tiny II}}$. As before, the basic idea is to let one or both of the $\lambda$ parameters approach $1$. There are two possible cases to consider: $\lambda_{1}\rightarrow 1$ at fixed $\lambda_{3}<1$, where only strings 1 and 2 (ending on $Q_{1}$ and $Q_{2}$) become infinitely long, and $\lambda_{1},\,\lambda_{3}\rightarrow 1$ , where all the strings become infinitely long.
We begin with the latter case, which leads to the $Y$-law, or more precisely to the term $\sigma L_{\text{min}}$ within it, as discussed in a3Q2008 . Our goal here is to determine the leading correction to this result. As the $\lambda$’s approach $1$, it follows immediately from (B.14) that $\cos\alpha_{1}=\cos\alpha_{3}=v\,\text{e}^{1-v}$. Substituting it into the first equation in (3.2) yields $\beta=\frac{\pi}{6}$. This implies that the point $Y$ coincides with the Fermat point of the triangle, as expected. It also implies that the cosine factor in (3.8) becomes $\cos\bigl{(}\tfrac{\pi}{6}+\tfrac{\theta}{2}\bigr{)}$. Since it must be positive, the apex angle is constrained by $\theta<\frac{2}{3}\pi$ [^6] For $\theta=\frac{2}{3}\pi$, this limit is undefined. It is straightforward to determine the value of $v$ corresponding to $\lambda_{1}=\lambda_{3}=1$. With $\alpha_{1}=\alpha_{3}$, the second equation in (3.2) gives $\sin\alpha_{1}=\mathsf{k}(1+4v)\text{e}^{-3v}$. Combining this with the expression for $\cos\alpha_{1}$, leads to the equation
@@
$\begin{split}E^{\,\text{\tiny II}}-2\cos(\tfrac{\theta}{2}-\tfrac{\pi}{3})\,\sigma\ell=&\mathsf{g}\sqrt{\mathsf{s}}\Bigl{(}2\bigl{(}{\cal E}^{-}(\lambda_{1},v)-\text{e}{\cal L}^{-}(\lambda_{1},v)\bigr{)}+{\cal E}^{-}(\lambda_{3},v)-\text{e}{\cal L}^{-}(\lambda_{3},v)+3\mathsf{k}\frac{\text{e}^{-2v}}{\sqrt{v}}\Bigr{)}+3c\\ &+2\bigl{(}1-\cos(\beta-\tfrac{\pi}{6})\bigr{)}\sin\tfrac{\theta}{2}\cos^{-1}\beta\,\sigma\ell\,.\end{split}$ (3.15)
@@
where the function ${\cal I}$ is defined in Appendix A. In the last step, we used the fact that the factor $1-\cos(\beta-\tfrac{\pi}{6})$ behaves like a power law in $1-\lambda_{1}$, whereas $\ell$ like a logarithm. This result extends the result of a3Q2016 from the equilateral triangle geometry to the isosceles one. The main conclusions then are that the leading correction is independent of $\theta$ and includes equal contributions from all the strings, enhanced by the contribution from the baryon vertex. Thus, in this limit, all the strings enter on equal footing.
@@
which shows explicit dependence on the apex angle $\theta$. Accordingly, the solution $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ is $\theta$-dependent and establishes an upper bound on $v$. The allowed range of $\theta$ is governed by the two conditions: $\lambda_{3}=1$ and $\alpha_{3}=0$, as discussed in Appendix F.
@@
To compute the constant term, we consider the difference between $E^{\,\text{\tiny II}}$ and the leading term appearing in the $-\ell$ expansion
@@
$-{E}^{\,\text{\tiny II}}=2\sigma\ell+\tilde{C}^{\,\text{\tiny II}}+o(1)\,,\qquad\text{with}\qquad\tilde{C}^{\,\text{\tiny II}}=3c-\mathsf{g}\sqrt{\mathsf{s}}\Bigl{(}2{\cal I}({\text{v}}_{\text{\tiny 1}}^{\text{\small-}})-{\cal E}^{-}(\lambda_{3},{\text{v}}_{\text{\tiny 1}}^{\text{\small-}})+2\text{e}\cos\tfrac{\theta}{2}{\cal L}^{-}(\lambda_{3},{\text{v}}_{\text{\tiny 1}}^{\text{\small-}})-3\mathsf{k}\frac{\text{e}^{-2{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}}}{\sqrt{{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}}}\Bigr{)}\,.$ (3.22)
@@ III.1.3 Configuration III
Finally, consider configuration $III$. A straightforward way to proceed is as follows. Since $\alpha_{3}$ changes its sign from negative to positive, the expressions for the geometrical constraint, leg length, and energy can be obtained from those of subsection 2 by respectively replacing ${\cal L}^{-}$ and ${\cal E}^{-}$ with ${\cal L}^{+}$ and ${\cal E}^{+}$. The force balance equations, however, remain unchanged. Thus, we can rewrite (3.8) and (3.9) as
@@
As before, when studying the $-\ell$ behavior, we take the limit $v\rightarrow 0$, in which the geometrical constraint becomes
@@
The solution ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ to this equation is $\theta$-dependent and establishes an upper bound on the parameter $v$. One might intuitively expect the range of $\theta$ to be set by the two conditions: $\alpha_{3}=0$ and $\alpha_{3}=\frac{\pi}{2}$. While this is true for the latter case, leading to $\theta=\pi$, it does not hold for the former. This is because ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is a double-valued function of $\theta$ [^8] For more on this, see Appendix D. Second, in this limit, strings 1 and 2 become infinitely long, while string 3 remains finite and contributes only to the subleading constant term in the expansion in $\ell$. This allows us to deduce directly from (3.22) that
@@ III.2 The simplest examples
[…]
@@ III.2.1 $\theta<\frac{2}{3}\pi$
Now consider $\theta=\frac{\pi}{6}$, a typical example with an acute apex angle. As explained in Appendix D, configuration $I$ dominates at small $\ell$, and is then replaced by configuration $II$ for larger values of $\ell$. Consequently, the potential is defined parametrically by two piecewise functions
=$\ell=\begin{cases}\ell^{\,\text{\tiny I}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny I}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}$ (3.30)
@@
=$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,.\end{cases}$ (3.32)
@@
$\sin\beta=\frac{1}{2}\cos^{-1}\alpha_{1}\,,\qquad\cos\alpha_{1}=\sqrt{1-\tfrac{9}{4}\mathsf{k}^{2}(1+4v)^{2}\text{e}^{-6v}}\,,\qquad\cos\bigl{(}\beta+\tfrac{\pi}{4}\bigr{)}{\cal L}^{-}(\lambda_{1},v)=\tfrac{1}{\sqrt{2}}\,{\cal L}^{+}(0,v)\,.$ (3.33)
@@
All of them agree with the $-\ell$ asymptotics
@@
As first noted in a3Q2016 , the constant terms appearing in the large- and $-\ell$ expansions are not equal. This feature sets the five-dimensional string model apart from phenomenological ansätze. It is noteworthy that the difference between these constants is well-defined (scheme-independent), and a straightforward estimate yields
@@
Thus, the constant term in the $-\ell$ expansion is smaller than that in the $-\ell$ expansion.
It is also instructive to make an estimate in relation to the $\Delta$-ansatz. Assuming that $E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is given by (C.1), the difference between the constant terms in the $-\ell$ expansions is estimated to be
@@
Now, let us present some estimates for the $-\ell$ behavior. The results are summarized in Table 1.

### v3

@@ II Preliminaries
Figure 1: A typical baryonic configuration in five dimensions. The heavy quarks are denoted by $Q_{i}$, the baryon vertex by $V$, its projection onto the boundary by $Y$, and the tangent angles of the strings by $\alpha_{i}$ ($i=1,\dots,3$). The arrows indicate the forces acting on the vertex, which is located in the bulk at $r=r_{v}$.
@@
$\mathbf{f}=\Bigl(0,\,0,\,-\tau_{v}\partial_{r_{v}}\,\frac{\text{e}^{-2\mathsf{s}r_{v}^{2}}}{r_{v}}\Bigr)\,.$ (2.6)
@@ III Isosceles Triangle Geometry
which are valid for $0\leq\beta\leq(\pi-\theta)/2$.
Because of axial symmetry, $\alpha_{1}=\alpha_{2}$, and the point $Y$ lies on the $y$-axis. In this case, the expressions for the string tension vectors take the form $\mathbf{e}_{1}=-\mathsf{g}w(r_{v})(\cos\beta\cos\alpha_{1},\sin\beta\cos\alpha_{1},\sin\alpha_{1})$, $\mathbf{e}_{2}=\mathsf{g}w(r_{v})(\cos\beta\cos\alpha_{1},-\sin\beta\cos\alpha_{1},-\sin\alpha_{1})$, and $\mathbf{e}_{3}=\mathsf{g}w(r_{v})(0,\cos\alpha_{3},-\sin\alpha_{3})$, with $w(r_{v})=\frac{\text{e}^{\mathsf{s}r_{v}^{2}}}{r_{v}^{2}}$. As a result, the force balance equation (2.5) reduces to two non-trivial components, corresponding to the $y$- and $r$-directions:
@@
Here $v=\mathsf{s}r_{v}^{2}$ and $\mathsf{k}=\frac{\tau_{v}}{3\mathsf{g}}$. It is important to note that for $\beta>(\pi-\theta)/2$, the string configuration does not exist, as the force balance equation has no solution in this case.
@@ III.1.1 Configuration I
[…]
@@
Here $\mathsf{g}=\frac{R^{2}}{2\pi\alpha^{\prime}}$ and $c$ is a normalization constant. Combined with the geometrical constraint and the force balance equations, this yields the function $E^{\,\text{\tiny I}}(\ell)$ in parametric form. [^5] The allowed range of $v$ is discussed in subsection B and Appendix D.
[…]
@@
$\cos\bigl(\beta+\tfrac{\theta}{2}\bigr)\sqrt{\cos\alpha_{3}}\,I(\cos^{2}\hskip-2.0pt\alpha_{1},\tfrac{3}{4},\tfrac{1}{2})=\sin\tfrac{\theta}{2}\sqrt{\cos\alpha_{1}}\bigl(1+I(\sin^{2}\hskip-2.0pt\alpha_{3},\tfrac{1}{2},\tfrac{3}{4})\bigr)\,$, (3.6)
@@ III.1.2 Configuration II
[…]
@@
It is now straightforward to analyze the small-$\ell$ behavior of $E^{\,\text{\tiny II}}(\ell)$. The force balance equations (3.5) apply in this case as well. A similar analysis to the one carried out previously shows that the geometrical constraint becomes
$\cos\bigl(\beta+\tfrac{\theta}{2}\bigr)\sqrt{\cos\alpha_{3}}\bigl(1+I(\sin^{2}\hskip-2.0pt\alpha_{1},\tfrac{1}{2},\tfrac{3}{4})\bigr)=\sin\tfrac{\theta}{2}\sqrt{\cos\alpha_{1}}\bigl(1+I(\sin^{2}\hskip-2.0pt\alpha_{3},\tfrac{1}{2},\tfrac{3}{4})\bigr)\,.$ (3.10)
@@
Now we would like to examine the large-$\ell$ behavior of $E^{\,\text{\tiny II}}$. As before, the basic idea is to let one or both of the $\lambda$ parameters approach $1$. There are two possible cases to consider: $\lambda_{1}\rightarrow 1$ at fixed $\lambda_{3}<1$, where only strings 1 and 2 (ending on $Q_{1}$ and $Q_{2}$) become infinitely long, and $\lambda_{1},\,\lambda_{3}\rightarrow 1$ , where all the strings become infinitely long.
We begin with the latter case, which leads to the $Y$-law, or more precisely to the term $\sigma L_{\text{min}}$ within it, as discussed in a3Q2008 . Our goal here is to determine the leading correction to this result. As the $\lambda$’s approach $1$, it follows immediately from (B.14) that $\cos\alpha_{1}=\cos\alpha_{3}=v\,\text{e}^{1-v}$. Substituting it into the first equation in (3.2) yields $\beta=\frac{\pi}{6}$. This implies that the point $Y$ coincides with the Fermat point of the triangle, as expected. It also implies that the cosine factor in (3.8) becomes $\cos\bigl(\tfrac{\pi}{6}+\tfrac{\theta}{2}\bigr)$. Since it must be positive, the apex angle is constrained by $\theta<\frac{2}{3}\pi$. [^6] For $\theta=\frac{2}{3}\pi$, this limit is undefined. It is straightforward to determine the value of $v$ corresponding to $\lambda_{1}=\lambda_{3}=1$. With $\alpha_{1}=\alpha_{3}$, the second equation in (3.2) gives $\sin\alpha_{1}=\mathsf{k}(1+4v)\text{e}^{-3v}$. Combining this with the expression for $\cos\alpha_{1}$, leads to the equation
@@
$\begin{split}E^{\,\text{\tiny II}}-2\cos(\tfrac{\theta}{2}-\tfrac{\pi}{3})\,\sigma\ell=&\mathsf{g}\sqrt{\mathsf{s}}\Bigl(2\bigl({\cal E}^{-}(\lambda_{1},v)-\text{e}{\cal L}^{-}(\lambda_{1},v)\bigr)+{\cal E}^{-}(\lambda_{3},v)-\text{e}{\cal L}^{-}(\lambda_{3},v)+3\mathsf{k}\frac{\text{e}^{-2v}}{\sqrt{v}}\Bigr)+3c\\ &+2\bigl(1-\cos(\beta-\tfrac{\pi}{6})\bigr)\sin\tfrac{\theta}{2}\cos^{-1}\hskip-3.0pt\beta\,\sigma\ell\,.\end{split}$ (3.15)
@@
[…]
@@
which shows explicit dependence on the apex angle $\theta$. Accordingly, the solution ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ is $\theta$-dependent and establishes an upper bound on $v$. The allowed range of $\theta$ is governed by the two conditions: $\lambda_{3}=1$ and $\alpha_{3}=0$, as discussed in Appendix F.
@@
To compute the constant term, we consider the difference between $E^{\,\text{\tiny II}}$ and the leading term appearing in the large-$\ell$ expansion
@@
${E}^{\,\text{\tiny II}}=2\sigma\ell+\tilde{C}^{\,\text{\tiny II}}+o(1)\,,\qquad\text{with}\qquad\tilde{C}^{\,\text{\tiny II}}=3c-\mathsf{g}\sqrt{\mathsf{s}}\Bigl(2{\cal I}({\text{v}}_{\text{\tiny 1}}^{\text{\small-}})-{\cal E}^{-}(\lambda_{3},{\text{v}}_{\text{\tiny 1}}^{\text{\small-}})+2\text{e}\cos\tfrac{\theta}{2}{\cal L}^{-}(\lambda_{3},{\text{v}}_{\text{\tiny 1}}^{\text{\small-}})-3\mathsf{k}\frac{\text{e}^{-2{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}}}{\sqrt{{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}}}\Bigr)\,.$ (3.22)
@@ III.1.3 Configuration III
[…]
@@
As before, when studying the small-$\ell$ behavior, we take the limit $v\rightarrow 0$, in which the geometrical constraint becomes
@@
The solution ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ to this equation is $\theta$-dependent and establishes an upper bound on the parameter $v$. One might intuitively expect the range of $\theta$ to be set by the two conditions: $\alpha_{3}=0$ and $\alpha_{3}=\frac{\pi}{2}$. While this is true for the latter case, leading to $\theta=\pi$, it does not hold for the former. This is because ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is a double-valued function of $\theta$. [^8] For more on this, see Appendix D. Second, in this limit, strings 1 and 2 become infinitely long, while string 3 remains finite and contributes only to the subleading constant term in the expansion in $\ell$. This allows us to deduce directly from (3.22) that
@@ III.2 The simplest examples
Having understood the basic string configurations, we can now proceed to compute the three-quark potential. But before doing so, we need to specify the model parameters. For the purposes of this paper, it is natural to use those from a3Q2016 , which were obtained by fitting lattice QCD data for pure $SU(3)$ gauge theory to the string model under consideration. Accordingly, we set $\mathsf{g}=0.176$, $\mathsf{s}=0.44\,\text{GeV}^{2}$, and $\mathsf{k}=-0.083$. [^9] Note that string configurations with all non-negative tangent angles are not allowed due to the negative value of $\mathsf{k}$. These values will be used in all subsequent estimates unless otherwise specified. Here, we do not attempt to compute the three-quark potential for all values of $\theta$. Instead, our goal is to illustrate the above ideas through a few examples, which we hope will clarify the general approach. A more complete analysis is presented in Appendix D.
@@ III.2.1 $\theta<\frac{2}{3}\pi$
Now consider $\theta=\frac{\pi}{6}$, a typical example with an acute apex angle. As explained in Appendix D, configuration II dominates at small $\ell$, and is then replaced by configuration $II$ for larger values of $\ell$. Consequently, the potential is defined parametrically by two piecewise functions
@@
where ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$ is the solution to the equation $\alpha_{1}(v)=0$. [^10] The upper index in ${\text{v}}_{\text{\tiny 0}}^{\text{\footnotesize i}}$ refers to a tangent angle, and the lower to its value. At this value of $v$ the tangent angle $\alpha_{1}$ changes the sign from positive to negative, signaling a transition between the string configurations. The value of ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$ can be found numerically from Eqs.(3.2) and (3.3) which take the form
@@
$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)&\text{for}\quad 0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,.\end{cases}$ (3.32)
@@
$\sin\beta=\frac{1}{2}\cos^{-1}\hskip-2.0pt\alpha_{1}\,,\qquad\cos\alpha_{1}=\sqrt{1-\tfrac{9}{4}\mathsf{k}^{2}(1+4v)^{2}\text{e}^{-6v}}\,,\qquad\cos\bigl(\beta+\tfrac{\pi}{4}\bigr){\cal L}^{-}(\lambda_{1},v)=\tfrac{1}{\sqrt{2}}\,{\cal L}^{+}(0,v)\,.$ (3.33)
@@
All of them agree with the large-$\ell$ asymptotics
@@
As first noted in a3Q2016 , the constant terms appearing in the large- and small-$\ell$ expansions are not equal. This feature sets the five-dimensional string model apart from phenomenological ansätze. It is noteworthy that the difference between these constants is well-defined (scheme-independent), and a straightforward estimate yields
@@
Thus, the constant term in the large-$\ell$ expansion is smaller than that in the small-$\ell$ expansion.
It is also instructive to make an estimate in relation to the $\Delta$-ansatz. Assuming that $E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is given by (C.1), the difference between the constant terms in the large-$\ell$ expansions is estimated to be
@@
Now, let us present some estimates for the small-$\ell$ behavior. The results are summarized in Table 1.